ABC Inc. 2018 sales are $1,100,000. Operating costs (excluding depreciation) are 70% of sales. Net fixed assets are $205,000. Depreciation amounted to 15% of net fixed assets. Interest expenses are $100,000. The tax bill must be calculated using the corporate income tax table in the text, and ABC Inc. paid 8% of net income in dividends.

ABC Income Statement

2018

Sales...............................................................                                                                                                                 

Operating costs (excluding depreciation).............                                            

EBITDA..............................................................                                                                                                           

Depreciation.......................................................                                                                                                  

EBIT ..................................................................                                                                                                          

Interest Expense.................................................                                                                                          

EBT....................................................................                                                                                                                 

Taxes*** ............. ..............................................                                                                                                     

Net Income........................................................                                                                                            

Common Dividends....................................                                                                                   

Addition to Retained Earnings ....................                                                               

Operating Cash Flow..................................

***For Tax Rate use the corporate tax table in chapter 2 to determine the company’s tax bill, average and marginal tax rates. Use the average tax rate for this portion of the income statement.

Average tax ________                                                                                

Marginal tax ________                

Corporate Tax Table

84 Part 1 The Company and Its Environment 2-9 The Federal Income Tax System The value of any financial asset (including stocks, bonds, and mortgages), as well as most real assets such as plants or even entire firms, depends on the after-tax stream of cash flows produced by the asset. The following sections describe the key features of corporate and individual taxation. 2-9a Corporate Income Taxes The corporate tax structure, shown in Table 2-1, is relatively simple. The marginal tax rate is the rate paid on the last dollar of income, while the average tax rate is the average rate paid on all income. To illustrate, if a firm had $65,000 of taxable income, its tax bill would be: Taxes$7,500 +0.25($65,000-$50,000) $7,500+S3,750 S11,250 Its marginal rate would be 25%, and its average tax rate would be $11,250/$65,000-17.3%. Note that corporate income above $18,333,333 has an average and marginal tax rate of 35%. INTEREST AND DIVIDEND INCOME RECEIVED BY A CORPORATION Interest income received by a corporation is taxed as ordinary income at regular corporate tax rates. However, 70% of the dividends received by one corporation from another are excluded from taxable income, while the remaining 30% are taxed at the ordinary tax rate.11 Thus, a corporation earning more than $18,333,333 and paying a 35% marginal tax rate would pay only (0.30)(0.35-0.105-10.5% of its dividend TABLE 2-1 Corporate Tax Rates as of January 2015 If a Corporations Taxable Income Is It Pays This Amount on the Base of the Bracket Plus This Percentage on the Excess over the Base Up to $50,000 $50,000-$75,000 $75,000-$100,000 $100,000-$335,000 335,000-$10,000,000 $10,000,000-$15,000,000 $15,000,000-$18,333,333 Over $18,333,333 $0 $7,500 $13,750 $22,250 $113,900 $3,400,000 $5,150,000 $6,416,667 15% 25 34 39 34 35 38 35 Average Tax Rate at Top of Bracket 15.0% 18.3 22.3 34.0 34.0 34.3 35.0 35.0 Source: See the IRS instructions for Form 1120 at www.irs.gov/pubfirs-pdf i1120.pdf IThe size of the dividend exclusion actually depends on the degree of ownership. Corporations that own less than 20% of the stock of the dividend-paying company can exclude 70% of the dividends received; firms that own more than 20% but less than 80% can exclude 80% of the dividends, and firms that ovvn more than 80% can exclude the entire dividend payment. we will, in general, assume a 70% dividend exclusion. Cepyrighi 2017 Cengage Learning. All Rights Reserved. May nx be copied scanned, onr depl caed in whele or in part. Due e eleconic rights, some third party contenn may be sappressed from the eBook asdfee eChaprens bi rial review has deened tar ay appressed c esen áes not materially affect he overall leaning experience. Cengage Learning Deserves te r too remove additional content at any dme if su bse ent n t restrict ons require it.
Answer #1

Net Fixed Assets = $ 205000

Sales = $ 1100000

LESS: Operating Costs (excluding depreciation) = 70 % of Sales = $ 770000

EBITDA = $ 330000

LESS: Depreciation = 15 % of Net Fixed Assets = 0.15 x 205000 = $ 30750

EBIT = $ 299250

LESS: Interest Expense = $ 100000

EBT = $ 199250

LESS: Taxes = 22250 + (199250 - 100000) x 0.39 = $ 22250 + 38707.5 = $ 60957.5

Net Income = $ 138292.5

Common Dividend = 8 % of Net Income = 0.08 x 138292.5 = $ 11063.4

Addition to Retained Earnings = (138292.5 - 11063.4) = $ 127229.1

Operating Cash Flow = Net Income + Depreciation = 138292.5 + 100000 = $ 238292.5

Average Tax Rate = Taxes / EBT = (60957.5 / 199250) = 0.3059 or 30.59 %

Marginal Tax Rate = Applicable Tax Bracket for the Firm = 39 %
Answer #2

How did you find the OCF?